# Bright and dark states of light: The quantum origin of classical interference

### v2

@@ Acknowledgements.
C.J.V.-B., C.E.M, and R.B. and are supported by the São Paulo Research Foundation (FAPESP, Grants Nos. 2017/13250-6, 2018/15554-5, 2019/11999-5, 2018/22402-7, 2023/01213-0, and 2022/00209-6) and by the Brazilian National Council for Scientific and Technological Development (CNPq, Grants Nos. 201765/2020-9, 402660/2019-6, 313886/2020-2, 409946/2018-4, 307077/2018-7, 465469/2014-0, and 311612/2021-0). C.E.M. acknowledges funding from the SNF, project number IZBRZ2-186312/1. We thank Markus Hennrich and Tobias Donner for comments. We also thank Wolfgang Schleich for valuable discussions concerning the formal description of the double-slit experiment. C.J.V.-B. and R.B. also thank the Max Planck Institute of Quantum Optics for their pleasant hospitality.

### v3

@@ Acknowledgements.
C.J.V.-B., C.E.M, and R.B., and are supported by the São Paulo Research Foundation (FAPESP, Grants Nos. 2017/13250-6, 2018/15554-5, 2019/11999-5, 2018/22402-7, 2023/01213-0, 2023/03300-7, and 2022/00209-6) and by the Brazilian National Council for Scientific and Technological Development (CNPq, Grants Nos. 201765/2020-9, 402660/2019-6, 313886/2020-2, 409946/2018-4, 307077/2018-7, 465469/2014-0, and 311612/2021-0). C.E.M. acknowledges funding from the SNF, project number IZBRZ2-186312/1. We thank Tobias Donner and Markus Hennrich for their comments and Wolfgang Schleich for discussions concerning the formal description of the double-slit experiment. C.J.V.-B. and R.B. thank the Max Planck Institute of Quantum Optics for their hospitality.